jabez177 (jabez177):

What is the solution to the equation 2(2 - j) = 4(j + 10) ?

OpenStudy (alexandervonhumboldt2):

4-2j=4j+40

OpenStudy (alexandervonhumboldt2):

-6j=36 j=?????????????

jabez177 (jabez177):

1 is my answer I got, Am I correct ?

OpenStudy (alexandervonhumboldt2):

no -6

OpenStudy (alexandervonhumboldt2):

-6*-6=36

OpenStudy (anonymous):

yes what @AlexandervonHumboldt2 has shown is correct! always work towards isolating the variable:)

jabez177 (jabez177):

Okay thank you I have one more

OpenStudy (anonymous):

YOU HAVE TO DISTUBE

OpenStudy (alexandervonhumboldt2):

hmm but i have a go sotty

OpenStudy (alexandervonhumboldt2):

@iheartfood can you help him i have to go

jabez177 (jabez177):

The product of 3 and a number is 12 more than the sum of the number and 26. What is the number

OpenStudy (anonymous):

3X+12=26

jabez177 (jabez177):

so 26 thnx

jabez177 (jabez177):

Wait I don't think it is 26 I really need the asnwer because i have to answer this and then leave to go to my dental appointment

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

To solve this problem, we must find where 3x - 12 = 26 + x Move all terms containing x to the left, all other terms to the right. Add '-1x' to each side of the equation. -12 + 3x + -1x = 26 + x + -1x Combine like terms: 3x + -1x = 2x -12 + 2x = 26 + x + -1x Combine like terms: x + -1x = 0 -12 + 2x = 26 + 0 -12 + 2x = 26 Add '12' to each side of the equation.-12 + 12 + 2x = 26 + 12 Combine like terms: -12 + 12 = 0 0 + 2x = 26 + 12 2x = 26 + 12 Combine like terms: 26 + 12 = 38 2x = 38 Divide each side by '2'. x = 19 Simplifying x = 19 Check: 3*19 = 57 - 12 = 45 26 + 19 = 45

jabez177 (jabez177):

I got 4

OpenStudy (anonymous):

@jabez177

OpenStudy (anonymous):

this is the answer

jabez177 (jabez177):

Okay um thank you

OpenStudy (anonymous):

sorry, i had to leave too but looks like you got help!